# Output Prediction of Quantum Circuits based on Graph Neural Networks

### v2

@@ 4.1 Evaluation methodology
where $\hat{y}_{i}^{\left(k\right)}$ is the output prediction associated with the ground truth target value $y_{i}^{\left(k\right)}$, $\bar{y}$ is the average of the target values, $N_{test}$ is the number of circuits in the test set, and $N$ is the number of outputs. It is worth emphasizing that $R^{2}$ quantifies accuracy in relation to the intrinsic variance of the test data. This metric is suitable for fair comparisons across circuits of different scales, as circuits of varying sizes may exhibit a tendency for output values to cluster around or near the mean to varying degrees.
Dataset Setup. For the noisy and noiseless simulator datasets, we randomly generate 20000 random circuits for each case. In addition to noise information derived from real quantum devices (IBM Perth, IBM Lagos, IBM Nairobi, and IBM Jakarta), we also define a noise model, referred to as Simulated Noise, in which gate errors, $T_{1}$, $T_{2}$, and readout errors are fixed at predefined values. During the extrapolation process for qubits, we randomly generate 40000 random circuits as the dataset.
@@ 4.2.1 Single-qubit expectation values
Furthermore, we compare our proposed method with an approach that uses Convolutional Neural Networks (CNNs) for single-qubit expectation values prediction [26]. The CNNs-based approach encodes quantum circuits into feature matrices using one-hot encoding, which are then fed into the networks. However, due to the requirement that CNNs need training data with identical feature dimensions, i.e., feature matrices of the same size, the CNN-based approach cannot leverage optimization strategies for randomly generated quantum circuits with $N$ qubits and depth $P$. Additionally, unlike the GNNs-based approach, which can incorporate device noise information into the feature vectors of nodes, the CNN-based approach lacks an effective mechanism to encode device noise into its feature matrices. Consequently, it is unable to utilize noise information to improve the accuracy of single-qubit expectation values prediction for noisy circuits. To evaluate and compare the performance of the two methods under noiseless conditions, we randomly generate 20000 quantum circuits using the gate set $S=\left\{T,H,CNOT\right\}$. The single-qubit expectation values of these circuits are predicted using both the GNNs-based and CNNs-based approaches and the coefficient of determination $R^{2}$ is calculated for each. The comparative results are shown in figure 8. From the comparison shown in figure 8, it is evident that the GNNs-based approach outperforms the CNNs-based approach, particularly as the circuit depth $P$ increases. This superior performance is primarily due to the fact that the topology of GNNs is derived directly from the connectivity of the quantum circuit, enabling it to better capture the interactions between quantum gates compared to the structure of CNNs. This advantage becomes increasingly pronounced as the circuit depth grows.
@@
First, we generate 20000 quantum circuits with a fixed number of qubits $N$ and circuit depths $P$ randomly distributed between 5 and 11, using the gate set $S=\left\{T,H,CNOT\right\}$, under both noiseless and noisy conditions. The single-qubit expectation values are predicted using GNNs. Among these, 100 circuits are selected as the test set to visualize scatter plots of the predicted versus actual values and to calculate the coefficient of determination $R^{2}$. The prediction results under noiseless conditions are illustrated in the left panel of figure 9, while those under noisy conditions, where the noise model is derived from IBM Perth, are presented in the right panel of figure 9. The red dashed line in both panels represents the $y=x$ line.
@@
$N$ Without Noise Simulated Noise IBM Perth IBM Lagos IBM Nairobi IBM Jakarta
@@ 4.2.2 Model extrapolation for GNNs
Since GNNs can take quantum circuits with varying numbers of qubits as input, we partially investigate this characteristic in Section 4.2.1. In this subsection, we further explore the extrapolation capability of GNNs, referring to their ability to predict the properties of quantum circuits with more qubits than those in the training dataset using the same network parameters. The single-qubit expectation values are set as the prediction targets for GNNs. For each simulation, we randomly generate 40000 circuits with depths ranging from 5 to 11 from the quantum gate set $S=\left\{T,H,CNOT\right\}$, serving as the dataset. During training, for each circuit, the ‘Gate Qubit’ part of the node feature vectors is randomly assigned the required qubits, ensuring that all qubits indicated by the ‘Gate Qubit’ part are trained. We simulate the extrapolation performance of GNNs trained on datasets of 3-qubit, 5-qubit, and 7-qubit circuits, respectively. The results comparing the extrapolation performance of these GNNs models to circuits with 7, 11, and 16 qubits under noiseless and Simulated Noise conditions are shown in the left panel of figure 11. The extrapolation performance of GNNs compared to CNNs under noiseless conditions is shown in the right panel of figure 11, and $\widetilde{N}$ represents the number of qubits in the circuits of the training set.
@@ 4.2.3 Two-qubit expectation values
In addition to predicting single-qubit expectation values, GNNs can also be trained to predict two-qubit expectation values. We focus on the first two qubits, corresponding to the rescaled expectation value $z_{ij}$ defined in equation (7). Notably, the same GNNs model can predict the two-qubit expectation value for any pair of qubits by swapping the target qubits with the first two qubits. We randomly generate 20,000 circuits with depths ranging from 5 to 11 as the dataset and used GNNs to predict two-qubit expectation values under both noiseless conditions and noisy conditions with noise derived from IBM Perth. The results, including the coefficient of determination $R^{2}$, are presented in figure 12.
@@ 4.3 Circuit performance comparison
To extend the applicability of GNNs in PQCs, which have wide applications in the NISQ era, we investigate the capability of GNNs to predict the overall properties of a quantum circuit, specifically the ground state energy of a 4-qubit hydrogen molecule $H_{2}$. Following the approach illustrated in figure 5, we construct PQCs using the gate set $S^{\ast}=\left\{R_{x}\left(\theta\right),R_{y}\left(\beta\right),CNOT\right\}$. Subsequently, the ground state energy of $H_{2}$ is calculated using the Variational Quantum Eigensolver (VQE) algorithm, during which the parameters in the quantum circuit are iteratively optimized until convergence. The computed ground state energy of $H_{2}$ is then treated as the prediction target for GNNs and used as a performance evaluation metric for the PQCs. Based on this metric, we implement two comparative schemes for circuits with different structures. The first approach, referred to as Indirect Comparison, evaluates the performance of two quantum circuits by predicting their respective metrics separately and comparing them. We also propose a new approach called Direct Comparison, wherein GNNs directly predict the probability that one PQC outperforms another. In this scheme, the input data consists of both circuits encoded as graph structures. To obtain the labels, Min-Max normalization is first applied to the ground state energies calculated for all PQCs. The normalized ground state energy values of the two circuits are then differenced, divided by 2, and shifted by 0.5 to generate the labels for the circuit pairs. In the simulation experiments, we generate datasets of 800, 1000, 5000, and 10000 noisy PQCs, with their corresponding ground state energies calculated via the VQE algorithm. Two GNNs are trained to perform the predictions required for the Indirect Comparison and Direct Comparison schemes, respectively. The prediction results are compared with the actual outcomes to evaluate the accuracy of the performance comparisons between different circuit structures. Detailed simulation results are presented in Table 4 and figure 13, where $N_{train}$ represents the amount of data used for training.
@@
From Table 4 and figure 13, we observe that the Direct Comparison scheme, which directly predicts the relative performance between two PQCs of different structures, significantly outperforms the Indirect Comparison scheme by an average of 36.2% on the same dataset. While the indirect comparison scheme can enhance prediction accuracy by increasing the data volume, this approach demands substantial computational resources, particularly when executing the VQE algorithm on noisy PQCs. Additionally, if one aims to obtain a complete performance ranking of all circuits, employing a Bubble Sort algorithm can achieve this objective. The Direct Comparison scheme, in addition to providing superior circuit performance predictions over the Indirect Comparison scheme, offers an alternative perspective for the design of GNNs circuit predictors. This not only broadens the application scope of GNNs as circuit performance predictors but also contributes to the advancement of efficient PQC evaluation strategies in quantum computing.
@@ 5 Conclusion
We investigate the application of Graph Neural Networks (GNNs) to predict the outputs of quantum circuits, which can be constructed using either non-parameterized or parameterized quantum gates. The predictions encompass single-qubit expectation values, two-qubit expectation values, and the overall property of a quantum circuit. In our simulations, the feature vectors incorporate device noise information, including gate errors, $T_{1}$, $T_{2}$, and readout errors. As a result, the prediction performance under noisy conditions is comparable to that under noiseless conditions. Furthermore, when compared to qubit-scalable CNNs-based method, the GNNs-based approach demonstrates superior prediction performance, particularly in scenarios involving scaling from few-qubit to many-qubit circuits. In simulations comparing the Direct Comparison and Indirect Comparison schemes, the Direct Comparison scheme exhibits a significant performance improvement, establishing a robust foundation for applying GNNs-based predictors in PQCs. In the future, we aim to extend the GNNs-based predictor to applications in Quantum Architecture Search (QAS) and quantum kernel design.

### v3

@@ 4.1 Evaluation methodology
where $\hat{y}_{i}^{\left(k\right)}$ is the output prediction associated with the ground truth target value $_{i}^{\left(k\right)}, y¯\bar{y}$ is the average of the target values, $N_{test}$ is the number of circuits in the test set, and $N$ is the number of outputs. It is worth emphasizing that $R^{2}$ quantifies accuracy in relation to the intrinsic variance of the test data. This metric is suitable for fair comparisons across circuits of different scales, as circuits of varying sizes may exhibit a tendency for output values to cluster around or near the mean to varying degrees.
Dataset Setup. For the noisy and noiseless simulator datasets, we randomly generate 20000 random circuits for each case. In addition to noise information derived from real quantum devices (IBM Perth, IBM Lagos, IBM Nairobi, and IBM Jakarta), we also define a noise model, referred to as Simulated Noise, in which gate errors, $T_{1}$, $T_{2}$, and readout errors are fixed at predefined values. The detailed parameters of the noise model are provided in A. During the extrapolation process for qubits, we randomly generate 40000 random circuits as the dataset.
@@ 4.2.1 Single-qubit expectation values
[…]
@@
First, we generate 20000 quantum circuits with a fixed number of qubits $N$ and circuit depths $P$ randomly distributed between 5 and 11, using the gate set $S=\left\{T,H,CNOT\right\}$, under both noiseless and noisy conditions. The single-qubit expectation values are predicted using GNNs. Among these, 100 circuits are selected as the test set to visualize scatter plots of the predicted versus actual values and to calculate the coefficient of determination $R^{2}$. The prediction results under noiseless conditions are illustrated in the figure 9(a), while those under noisy conditions, where the noise model is derived from IBM Perth, are presented in the figure 9(b). The red dashed line in both panels represents the $y=x$ line.
@@
NN Without Noise Simulated Noise IBM Perth IBM Lagos IBM Nairobi IBM Jakarta
@@ 4.2.2 Model extrapolation for GNNs
Since GNNs can take quantum circuits with varying numbers of qubits as input, we partially investigate this characteristic in Section 4.2.1. In this subsection, we further explore the extrapolation capability of GNNs, referring to their ability to predict the properties of quantum circuits with more qubits than those in the training dataset using the same network parameters. The single-qubit expectation values are set as the prediction targets for GNNs. For each simulation, we randomly generate 40000 circuits with depths ranging from 5 to 11 from the quantum gate set $S=\left\{T,H,CNOT\right\}$, serving as the dataset. During training, for each circuit, the ‘Gate Qubit’ part of the node feature vectors is randomly assigned the required qubits, ensuring that all qubits indicated by the ‘Gate Qubit’ part are trained. We simulate the extrapolation performance of GNNs trained on datasets of 3-qubit, 5-qubit, and 7-qubit circuits, respectively. The results comparing the extrapolation performance of these GNNs models to circuits with 7, 11, and 16 qubits under noiseless and Simulated Noise conditions are shown in the figure 11(a). The extrapolation performance of GNNs compared to CNNs under noiseless conditions is shown in the figure 11(b), and $\widetilde{N}$ represents the number of qubits in the circuits of the training set.
@@ 4.2.3 Two-qubit expectation values
In addition to predicting single-qubit expectation values, GNNs can also be trained to predict two-qubit expectation values. We focus on the first two qubits, corresponding to the rescaled expectation value $z_{ij}$ defined in equation (8). Notably, the same GNNs model can predict the two-qubit expectation value for any pair of qubits by swapping the target qubits with the first two qubits. We randomly generate 20000 circuits with depths ranging from 5 to 11 as the dataset and used GNNs to predict two-qubit expectation values under both noiseless conditions and noisy conditions with noise derived from IBM Perth. The results, including the coefficient of determination $R^{2}$, are presented in figure 12.
Figure 12: Prediction of two-qubit expectation values with circuit depths of 5-11 under different conditions.
@@ 4.3 Circuit performance comparison
To extend the applicability of GNNs in PQCs, which have wide applications in the NISQ era, we investigate the capability of GNNs to predict the overall properties of a quantum circuit, specifically the ground state energy of a 4-qubit hydrogen molecule H2. Following the approach illustrated in figure 5, we construct PQCs using the gate set $S^{\ast}=\left\{R_{x}\left(\theta\right),R_{y}\left(\beta\right),CNOT\right\}$. Subsequently, the ground state energy of H2 is calculated using the Variational Quantum Eigensolver (VQE) algorithm, during which the parameters in the quantum circuit are iteratively optimized until convergence. The computed ground state energy of H2 is then treated as the prediction target for GNNs and used as a performance evaluation metric for the PQCs. Based on this metric, we implement two comparative schemes for circuits with different structures. The first approach, referred to as Indirect Comparison, evaluates the performance of two quantum circuits by predicting their respective metrics separately and comparing them. We also propose a new approach called Direct Comparison, wherein GNNs directly predict the probability that one PQC outperforms another. In this scheme, the input data consists of both circuits encoded as graph structures. To obtain the labels, Min-Max normalization is first applied to the ground state energies calculated for all PQCs. The normalized ground state energy values of the two circuits are then differenced, divided by 2, and shifted by 0.5 to generate the labels for the circuit pairs. In the simulation experiments, we generate datasets of 800, 1000, 5000, and 10000 noisy PQCs, with their corresponding ground state energies calculated via the VQE algorithm. Two GNNs are trained to perform the predictions required for the Indirect Comparison and Direct Comparison schemes, respectively.
Before comparing the prediction accuracy between the Direct and Indirect Comparison schemes, we first evaluate the performance of the same set of quantum circuits through 100 independent VQE experiments. Specifically, three evaluation methods are considered: (1) Direct Comparison, where GNNs predict the probability that one PQC outperforms another; (2) Indirect Comparison, where the performance of two circuits is compared based on their predicted metrics; and (3) Calculation Comparison, where both circuits are directly optimized using the VQE algorithm with the COBYLA optimizer for 200 iterations to obtain their ground-state energies. The mean and standard deviation of the results over 100 trials are summarized in Table 4 and illustrated in figure 13, where the vertical axis is plotted on a logarithmic scale. As shown in Table 4, the Direct Comparison and Indirect Comparison schemes using GNNs as predictors achieve $11321\times$ and $5941\times$ speedups, respectively, compared to the Calculation Comparison method based on direct VQE computation. These results demonstrate the potential of the trained GNNs to perform large-scale and rapid performance evaluation of PQCs, providing an efficient alternative to costly VQE simulations.
Table 4: Runtime of Direct Comparison, Indirect Comparison and Calculation Comparison
Scheme Direct Comparison Indirect Comparison Calculation Comparison
Runtime $\left(\cdot s\right)$ $00053\pm 0.0001 0.0$$00101\pm 0.0002$ $6.00012\pm 0.8178$
Speedup $\mathbf{11321\times}$ $\mathbf{5941\times}$ $\mathbf{1\times}$
Figure 13: Comparison of runtime of Direct Comparison, Indirect Comparison and Calculation Comparison.
Having established the runtime advantage of the GNN-based approaches, we next assess their prediction accuracy. The prediction results are compared with the actual outcomes to evaluate the accuracy of the performance comparisons between different circuit structures. Detailed simulation results are presented in Table 5 and figure 14, where $N_{train}$ represents the amount of data used for training.
@@
From Table 5 and figure 14, we observe that the Direct Comparison scheme, which directly predicts the relative performance between two PQCs of different structures, significantly outperforms the Indirect Comparison scheme by an average of 36.2% on the same dataset. While the indirect comparison scheme can enhance prediction accuracy by increasing the data volume, this approach demands substantial computational resources, particularly when executing the VQE algorithm on noisy PQCs. Additionally, if one aims to obtain a complete performance ranking of all circuits, employing a Bubble Sort algorithm can achieve this objective. The Direct Comparison scheme, in addition to providing superior circuit performance predictions over the Indirect Comparison scheme, offers an alternative perspective for the design of GNNs circuit predictors. For example, in the design of parameterized quantum circuits (PQCs), when facing a large number of candidate circuits to be screened and some high-performing reference circuits are already available, the Direct Comparison scheme can leverage these benchmark circuits to rapidly eliminate a large number of mediocre candidates, thereby reducing the cost of individually evaluating each parameterized circuit. This not only broadens the application scope of GNNs as circuit performance predictors but also contributes to the advancement of efficient PQC evaluation strategies in quantum computing.
@@ 5 Conclusion
We investigate the application of Graph Neural Networks (GNNs) to predict the outputs of quantum circuits, which can be constructed using either non-parameterized or parameterized quantum gates. The predictions encompass single-qubit expectation values, two-qubit expectation values, and the overall property of a quantum circuit. In our simulations, the feature vectors incorporate device noise information, including gate errors, $T_{1}$, $T_{2}$, and readout errors. As a result, the prediction performance under noisy conditions is comparable to that under noiseless conditions. Furthermore, when compared to qubit-scalable CNNs-based method, the GNNs-based approach demonstrates superior prediction performance, particularly in scenarios involving scaling from few-qubit to many-qubit circuits. In simulations comparing the Direct Comparison and Indirect Comparison schemes, the Direct Comparison scheme exhibits a significant performance improvement, establishing a robust foundation for applying GNNs-based predictors in PQCs. In the future, we aim to extend the GNNs-based predictor to predict other circuit properties using the same framework and apply it to quantum qrchitecture search (QAS) and quantum kernel design (QKD). While our simulations indicate that the predictor possesses a certain degree of extrapolation ability, how to further improve this ability is left for future investigation.
@@
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62471126), the Jiangsu Frontier Technology Research and Development Plan (Grant No. BF2025066), the Fundamental Research Funds for the Central Universities (Grant No. 2242022k60001), the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No.2022ZB139), and the Natural Science Foundation of the Higher Education Institutions of Jiangsu Province (Grant No.102024097).
Appendix A Parameters of the noise model
The noise information used in our experiments for the five noise models (IBM Perth, IBM Lagos, IBM Nairobi, IBM Jakarta, and the Customized Simulated Noise) is summarized in Tables 610. In real quantum devices, the supported gate sets vary, and different quantum gates may exhibit distinct noise characteristics depending on the qubits on which they are applied. For experimental convenience and without affecting the validity of the results, we categorize the error into single-qubit gate error and two-qubit gate error, while neglecting the effects of the actual device topology. In addition, in our Customized Simulated Noise model, all qubits share identical parameter values.
Table 6: The Noise Parameters of the IBM Perth device
Qubit $T_{1}$ $\left(\cdot us\right)$ $T_{2}$ $\left(\cdot us\right)$ Single-qubit Gate Error Two-qubit Gate Error Readout Error 0-1 Readout Error 1-0
1 197.79 97.02 0.0001836 0.00735 0.0310 0.0226
2 158.07 48.06 0.0003291 0.00770 0.0222 0.0236
3 278.36 83.10 0.0002153 0.00694 0.0282 0.0230
4 223.29 211.00 0.0002381 0.00835 0.0180 0.0130
5 109.91 126.96 0.0003141 0.01169 0.0230 0.0164
6 236.63 188.44 0.0002563 0.00981 0.0264 0.0208
7 193.96 291.83 0.0003357 0.00984 0.0072 0.0054
Table 7: The Noise Parameters of the IBM Lagos device
Qubit $T_{1}$ $\left(\cdot us\right)$ $T_{2}$ $\left(\cdot us\right)$ Single-qubit Gate Error Two-qubit Gate Error Readout Error 0-1 Readout Error 1-0
1 96.00 40.53 0.0001813 0.00655 0.0180 0.0092
2 131.71 75.18 0.0001439 0.00770 0.0192 0.0178
3 174.23 160.96 0.0002279 0.00685 0.0094 0.0074
4 139.45 84.68 0.0001988 0.00651 0.0166 0.0144
5 113.32 29.50 0.0002090 0.00723 0.0210 0.0198
6 131.19 64.52 0.0001975 0.00685 0.0148 0.0190
7 93.17 81.31 0.0001845 0.00588 0.0130 0.0152
Table 8: The Noise Parameters of the IBM Nairobi device
Qubit $T_{1}$ $\left(\cdot us\right)$ $T_{2}$ $\left(\cdot us\right)$ Single-qubit Gate Error Two-qubit Gate Error Readout Error 0-1 Readout Error 1-0
1 104.92 26.07 0.0002782 0.01034 0.0318 0.0110
2 141.98 97.62 0.0005325 0.01765 0.0720 0.0198
3 57.67 69.22 0.0036500 0.03446 0.0406 0.0116
4 145.77 59.13 0.0003363 0.01259 0.0366 0.0096
5 99.10 61.06 0.0002725 0.00954 0.0326 0.0098
6 126.61 16.25 0.0002855 0.01118 0.0432 0.0204
7 148.77 105.21 0.0002028 0.00696 0.0410 0.0142
Table 9: The Noise Parameters of the IBM Jakarta device
Qubit $T_{1}$ $\left(\cdot us\right)$ $T_{2}$ $\left(\cdot us\right)$ Single-qubit Gate Error Two-qubit Gate Error Readout Error 0-1 Readout Error 1-0
1 145.55 47.91 0.0003194 0.00806 0.0282 0.0070
2 143.21 28.18 0.0002136 0.00872 0.0396 0.0096
3 110.89 22.67 0.0002076 0.00914 0.0252 0.0082
4 78.53 36.52 0.0002149 0.00776 0.0254 0.0112
5 145.55 47.91 0.0003194 0.00806 0.0282 0.0070
6 128.70 65.54 0.0002627 0.00616 0.0646 0.0480
7 139.55 21.58 0.0002507 0.00577 0.0400 0.0244
Table 10: The Noise Parameters of Customized Simulated Noise
Qubit $T_{1}$ $\left(\cdot us\right)$ $T_{2}$ $\left(\cdot us\right)$ Single-qubit Gate Error Two-qubit Gate Error Readout Error 0-1 Readout Error 1-0
1 197.79 97.02 0.0001 0.0083 0.01 0.01
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
16 197.79 97.02 0.0001 0.0083 0.01 0.01